N arithmetic mean are inserted between the number 7 and 49 if the sum of these means is 364 then the sum of their squares is How to calculate the sum of their squares.

N arithmetic  mean are inserted between the number 7 and 49 if the sum of these means is 364 then the sum of their squares is
How to calculate the sum of their squares.

Grade:11

1 Answers

Arun
5 years ago
 
 
Dear student
 
a.m. between 7 and 49 = 7 + d, 7 + 2d, …....., 7+ nd
 
now
 
n/2 [14 + (n+ 1) d]  = 364
 
also
 
49 = 7 + (n+ 1) d
 
now there are 2 equations and 2 variables n and d now you can find